OpenStudy (anonymous):

Note: For questions 7 and 8, remember to show all the steps that you use to solve the problem. You can use the comments field to explain your work. Your teacher will reivew each step of your response to ensure you recieve proper credit for your answer. Use the function rule to complete the table. –10x + y = 4

OpenStudy (anonymous):



Directrix (directrix):

–10x + y = 4 ------------ x = -2. Substitute (-2) in for x in the equation –10x + y = 4. –10*(-2) + y = 4 20 + y = 4 y = 4 - 20 = -16 (-2 , -15) is the first entry in the chart. Follow the same pattern for x = -1.